OpenStudy (anonymous):

Solve: 38=4w-26 Work out

OpenStudy (jit4won):

38 =4w -26 38 + 26 = 4w -26 +26 64 = 4w 64/4 =4w/4 16= w

OpenStudy (jit4won):

any thing else ?

OpenStudy (anonymous):

ya\

OpenStudy (anonymous):

ill post it

OpenStudy (jit4won):

Hahaha

OpenStudy (jit4won):

Next question